OpenStudy (anonymous):

ind int ln(2x+1)dx=xln(2x+1)-x+(1/2)ln(2x+1)+c

OpenStudy (anonymous):

how do you get a final ans of (1/2)(2x+1)ln(2x+1)-x+c?

OpenStudy (anonymous):

how did you get the first answer?

OpenStudy (anonymous):

should be by parts, although it comes up so often it is easier just to remember that \[\int \ln(u)du=u\ln(u)-u+C\]

OpenStudy (anonymous):

too long to type out, I tried before....but I can't understand how to get from my answer to the final ans.

OpenStudy (anonymous):

ok let me see what we get using \(u=2x+1\) and \(\frac{1}{2}du =dx\)

OpenStudy (anonymous):

we can go right to the answer and get \[\frac{1}{2}\left(u\ln(u)-u\right)\]

OpenStudy (anonymous):

\[\frac{1}{2}\left((2x+1)\ln(2x+1)-(2x+1)\right)\]

OpenStudy (anonymous):

then multiply the second term by \(\frac{1}{2}\) to get \[\frac{1}{2}(2x+1)\ln(2x+1)-x-\frac{1}{2}\]

OpenStudy (anonymous):

oh i see what was bothering you the \(\frac{1}{2}\) at the end is unimportant, it is a constant

OpenStudy (anonymous):

it gets sucked up in to the \(+C\)

OpenStudy (anonymous):

i mean i guess that is what was bothering you, because if you do the algebra you do not get exactly what is written, but you are off by a constant which you always are